# Pricing AI Model Accuracy

### v1

@@ 3.3 Combination of Positively and Negatively Correlated Sensitivities
This trivially implies that prices and revenue are highest when consumers all have perfectly negatively correlated cost sensitivities, i.e., when $\zeta=0$. This result stems from the fact that when $\alpha=\beta$, consumers exhibit no preferential distinction between the FP and FN rates themselves. The market share for each firm remains significant even when prices are pushed up, so revenue increases.
@@ 4 Firm Flexibility in FP and FN Rates
In Section 3, we assume that firms are endowed with fixed FP and FN rates and have no ability to change them. We now introduce flexibility on the error dimension through investment; firms can invest to reduce their FP and FN rates firms. We consider such investments for both firms under both the strict and split domination cases and analyze the existence and direction of profitable deviations.
@@ 4.1 Investments for Inferior Firm in Strict Domination
Since firm 2 initially does not enter the market, they have zero revenue. Therefore, an investment that captures a strictly positive proportion of the market share is profitable for firm 2. We show that firm 2 can make strictly positive revenue by investing on its FP and FN rates; however, the sizes of such beneficial investments vary by dimension and the error rate differences. The distinction between investments in FP and FN rates stems from the volatility of error rates.
@@
Assume $|FN_{2}-FN_{1}|<|FP_{2}-FP_{1}|$, and firm 1 strictly dominates firm 2 on both dimensions. An investment by firm 2 which lowers $FN_{2}$ to $FN_{2}$ results in strictly positive revenue iff the investment achieves FN domination: i.e. $FN_{2}$ $<FN_{1}$.
@@ 4.2 Investments Under Split Domination
The theorem follows trivially from Lemmas 1 and 2. This result is slightly counter-intuitive. Firms should invest on further establishing their competitive advantage rather than trying to close the gap on the other dimension. Such investment behavior implies a degree of specialization in error rates across firms. If a firm is trying to produce a model exclusively for image classification in computer vision, it may not be very worried if the model is very inaccurate on classifying images of animals. Past research has touched on this as well.
@@
In addition, investing on their inferior dimension is equivalent to reducing the product differentiation in the market. This would imply that the firms would no longer be able to appeal as much to extreme consumers, resulting in lower and more similar prices between firms.
@@ 5 Welfare Analysis
Strictly profitable firm investments have a negative impact on consumer welfare and a strictly positive effect on firm welfare. The overall effect on welfare is positive (i.e., firm welfare + consumer welfare $>$ 0).
@@
[…]
where:
@@
Firm investments on their inferior dimension have a strictly positive effect on consumers and overall welfare.
For example, an investment by firm 1 towards lowering its FP rate (its inferior dimension) would result in strictly lower profit as shown by Theorem 4. Furthermore, firm 2 would experience a relatively larger drop in profit, due to losing market share but also having to reduce its price. However, consumers would be strictly better off due to lower prices. An interesting consequence of this is that firms can be incentivized to invest on their inferior dimension, which would be equivalent to reducing product differentiation in the market.
@@ 6 Summary and Conclusions
This paper develops a dynamic consumer-firm duopoly model to analyze optimal firm behavior through specialized investments. Each firm aims to minimize its model’s error, but this choice can often be suboptimal when firms compete and consumers have heterogeneous preferences. By decomposing model error into false positive and false negative rates, we allow firms to reduce error on one dimension at a time through investments.
@@
Our model assumes all information about consumer types is public information; introducing an element of private information into the consumer types could be interesting. Each consumer would be endowed with their own personal FP/FN rates and only buy a product from firms if the transaction is profitable error-wise. Our model assumes that every consumer must not only purchase from a firm, but we can also deterministically pinpoint from which firm they will buy from. Implicitly, we assume a mature and entirely covered market, in which we not only have a fixed market size with no growth, but all firms in the market are actually selling a product. A reasonable relaxation of the model could consider the possibility of shut downs when their product is no longer profitable, and similarly the potential for market entry if a firm comes up with a innovative product.
Another interesting extension would be introducing dynamics into the model and analyzing convergence rates to equilibria. The market clearing procedure could be repeated for multiple periods, where firms can change their investments and consumers can make different purchases period by period. This would relax the implicit consumer captivity assumption of the Hotelling model and allow for consumers to move around rather than committing to a side/firm.
@@ A.1 Equilibrium Prices in Perfectly Positively Correlated Model: $\alpha=\beta$
$\tilde{\alpha}=\begin{cases}0&\text{if}\ \frac{p_{2}-p_{1}}{\mathcal{F}_{1}-\mathcal{F}_{2}}\leq 0\\ \frac{p_{2}-p_{1}}{\mathcal{F}_{1}-\mathcal{F}_{2}}&\text{if}\ 0<\frac{p_{2}-p_{1}}{\mathcal{F}_{1}-\mathcal{F}_{2}}<1\\ 1&\text{if}\ \frac{p_{2}-p_{1}}{\mathcal{F}_{1}-\mathcal{F}_{2}}\geq 1.\\ \end{cases}$
@@ A.2 Equilibrium Prices with Split Domination in Perfectly Negatively Correlated Model: $\alpha=1-\beta$
$\tilde{\alpha}=\begin{cases}0&p_{2}-p_{1}\leq FN_{1}-FN_{2}\\ 1&p_{2}-p_{1}\geq FP_{1}-FP_{2}\\ \frac{p_{2}-p_{1}+FN_{2}-FN_{1}}{\Delta F_{1}-\Delta F_{2}}&\text{otherwise.}\\ \end{cases}$
@@
The firms’ reaction functions can then be solved using casework on $\tilde{\alpha}$ and relevant FOCs. Note that if $\tilde{\alpha}=0$, then $R_{1}=0,R_{2}=p_{2}$. Since firms are not subject to direct costs in this model, firm 2 will set price $p_{2}=\bar{p}$ and $p_{1}=0$. Similarly, if $\tilde{\alpha}=1$, firm 2’s revenue will be 0 and equilibrium prices will be $p_{1}=\bar{p}$ and $p_{2}=0$. When $10<\tilde{\alpha}<$ 1, the reaction functions can be calculated by taking the FOCs of firm revenues with respect to prices.
@@ A.3 Equilibrium Prices with Strict Domination in Perfectly Negatively Correlated Model: $\alpha=1-\beta$
$\tilde{\alpha}=\begin{cases}0&p_{2}-p_{1}>FN_{1}-FN_{2}\\ 1&p_{2}-p_{1}<FP_{1}-FP_{2}\\ \frac{p_{2}-p_{1}+FN_{2}-FN_{1}}{\Delta F_{1}-\Delta F_{2}}&\text{otherwise.}\\ \end{cases}$ (2)

### v2

@@ 3.3 Combination of Positively and Negatively Correlated Sensitivities
This trivially implies that prices and revenue are highest when consumers all have perfectly negatively correlated cost sensitivities, i.e., when $\zeta=0$. See Appendix C for the relevant graph of firm prices and revenues across varying $\zeta$. This result stems from the fact that when $\alpha=\beta$, consumers exhibit no preferential distinction between the FP and FN rates themselves. The market share for each firm remains significant even when prices are pushed up, so revenue increases.
@@ 4 Firm Flexibility in FP and FN Rates
In Section 3, we assume that firms are endowed with fixed FP and FN rates and have no ability to change them. We now introduce flexibility on the error dimension through investment; firms can invest to reduce their FP and FN rates. We consider such investments for both firms under both the strict and split domination cases and analyze the existence and direction of profitable deviations.
@@ 4.1 Investments for Inferior Firm in Strict Domination
Since firm 2 initially does not enter the market, they have zero revenue. Therefore, an investment that captures a strictly positive proportion of the market share is profitable for firm 2. We show that firm 2 can make strictly positive revenue by investing on its FP and FN rates; however, the sizes of such beneficial investments vary by dimension and error rate differences. The distinction between investments in FP and FN rates stems from the volatility of error rates.
@@
Assume $|FN_{2}-FN_{1}|<|FP_{2}-FP_{1}|$, and firm 1 strictly dominates firm 2 on both dimensions. An investment by firm 2 which lowers $FN_{2}$ to $FN_{2}$ results in strictly positive revenue iff the investment achieves FN domination: i.e. $FN_{2}$ $FN_{1}$.
@@ 4.2 Investments Under Split Domination
The theorem follows trivially from Lemmas 1 and 2. This result is slightly counter-intuitive. Firms should invest on further establishing their competitive advantage rather than trying to close the gap on the other dimension. Such investment behavior implies a degree of specialization in error rates across firms. If a firm is trying to produce a model exclusively for image classification in computer vision, it may not be very worried if the model is very inaccurate on classifying images of animals. Past research has also touched on this point.
@@
In addition, investing on their inferior dimension is equivalent to reducing the product differentiation in the market. This would imply that firms would no longer be able to appeal as much to extreme consumers, resulting in lower and more similar prices between firms.
4.3 Investment Costs
We now consider the case with an explicit investment cost function for each firm. In addition to investing on reducing their error rates explicitly, we introduce a latent variable for the size of each firm’s training dataset denoted as $S$. Firms are able to invest in two ways: increasing the size of their training dataset and explicitly improving their error rates. A size investment will have more indirect benefits to the firm: any future investments on error rates will be less costly.
Much of the artificial intelligence and machine learning literature on scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022) has demonstrated that dataset size and performance/error are highly intertwined. In our framework, this implies that the effectiveness of firm investments are directly related to their latent dataset size variable. We propose that the training dataset size each firm has directly effects its investment costs: a firm with a much larger training dataset will be able to invest more efficiently on improving their error rates as opposed to a firm with much fewer training samples.
Marginal costs should also be different across FP and FN rates. There should a distinct tradeoff for each firm on reducing their FP and FN rates (Tirole, 1988; Mas-Colell et al., 1995). In particular, if a firm already has a low FP error rate, it should be more expensive to lower it even further. A firm with $FP=0.1$ and investing to reduce it to 0.09 should experience much higher costs than a firm attempting to lower its FP rate from 0.3 to 0.29.
Therefore, each firm’s investment cost function has the following general form:
$C(FP,FN,S)=\frac{g(S)}{2}\left[s_{\overline{FP}}(FP-\overline{FP})^{2}+s_{\overline{FN}}(FN-\overline{FN})^{2}\right]+f(S)$
where $f(S)=\lambda_{linear}(S-\bar{S})+\frac{\lambda_{convex}}{p}(S-\bar{S})^{p}+\lambda_{runtime}(S-\bar{S})$ (with $p\geq 2$), is an increasing convex function capturing the firm’s cost for increasing their training dataset size from $\bar{S}$ to $S$, and $g(S):S\rightarrow\mathbb{R}$ is a monotonically decreasing function that maps the firm’s choice of size to a cost scaling parameter.
Here, the bar variables $\overline{FP},\overline{FN},$ and $\overline{S}$ denote the firm’s current error rates and size, respectively. [^2] For notational clarity, we omit the firm-specific subscript $j$ for each error rate: all variables above are unique to each firm. The factor of 2 is included purely for cleanliness when differentiating. The $\lambda$ parameters specify how sensitive the firm is to different sources of scaling costs: $\lambda_{linear}$ specifies linear scaling costs for increasing the number of training samples, $\lambda_{convex}$ specifies convex scaling costs (such as potentially regulatory or operational costs), and $\lambda_{runtime}$ specifies the additional degree of costs incurred from greater runtime during training. Finally, $s_{\overline{FP}}$ and $s_{\overline{FN}}$ capture the fact that error rates are more costly to lower when they are already low; a simple specification could set $s_{\overline{FP}}=\frac{1}{\overline{FP}}$, $s_{\overline{FN}}=\frac{1}{\overline{FN}}$. Note that $C(\overline{FP},\overline{FN},\overline{S})=0$.
We first discuss the case of split domination. Under the results from Section 4.2, neither firm will have an incentive to invest on improving the error rates on its strictly inferior dimension. Therefore, the addition of a non-zero cost function will only dis-incentivize such investments; we implicitly focus on the case of investments on one’s superior dimension.
Under this structure, firms will invest on two dimensions: their superior dimension and firm size. Although size directly affects the firm’s cost for improving their error rates, it acts as a latent variable in the firm’s analysis. The firm has some size level that is private information; it then chooses an optimal level of investment on both size and error rates to improve $\overline{FP}$ and $\overline{FN}$ to $FP$ and $FN$ respectively. Then, these error rates are reported to all consumers, and price equilibria are determined through market clearing. Formally, the firm’s revenue is unaffected by their training dataset size. This resembles real-world scenarios in which a larger training dataset may allow for more efficient model investments, but in the end, the models themselves and their attractiveness to consumers are not directly related to the firm’s training dataset size.
We can thus compute the firm’s optimal level of investment on both such dimensions using a standard optimization argument. Firms will choose investment levels such that the marginal cost and marginal revenue are equal, and we note from above that firm 1 dominates on the FN dimension and firm 2 on the FP dimension. For simplicity, we focus on the analysis from firm 1’s perspective; the analysis is symmetric for firm 2.
Corollary 1 (Optimal Choices Under AI-Based Cost Functions for Firm 1).
Firm 1’s investment choices $(FN_{1}^{*},S_{1}^{*})$ are an interior optimum and satisfy the following:
$\frac{1}{2}g^{\prime}(S_{1}^{*})\big{[}s_{\overline{FP_{1}}}(FP_{1}-\overline{FP_{1}})^{2}+s_{\overline{FN_{1}}}(FN_{1}-\overline{FN_{1}})^{2}\big{]}+f^{\prime}(S_{1}^{*})=0$
and
$FN_{1}^{*}=\overline{FN_{1}}+\frac{\partial R_{1}/\partial FN_{1}}{g(S_{1}^{*})s_{\overline{FN_{1}}}}$
The first condition presents an optimality condition on $S_{1}^{*}$: since revenue is not directly affected by the firm’s latent size variable, this is just an optimization condition on the cost function. Note that since $g(S)$ is decreasing and $f(S)$ is convex, such an $S_{1}^{*}$ exists. The second condition is a more standard firm profit maximization condition; details are provided in Appendix B.3.
The second condition implies that the size of the firm’s optimal investment is strongly dependent on its scalability and growth potential captured by $g(S)$. In particular, if a firm is able to make use of its training data very efficiently (i.e., having a lower $g(S)$), then the optimal error rate it should aim for ($FN_{1}^{*}$ in the case above) is much higher.
We can do a similar analysis under the strict domination case. As discussed in Proposition 2, the entrant can invest on either dimension, and it chooses the optimal investment decision depending on the ordering between the firms’ FP and FN rates. Since now a lump-sum investment by the entrant on either dimension leads to successful entry, we consider investments on both dimensions. However, unless the entrant faces a very forgiving cost function (that is, $g(S)$ is very small), entry is very difficult to incentivize. Even if the entrant does successfully obtain a non-zero market share, it will receive a very small level of revenue and is unlikely to have this offset the incurred costs.
Building on Section 4.1, when firm 1 strictly dominates firm 2, a profitable investment tuple $(FP_{2},FN_{2},S_{2})$ by firm 2 must change the ordering of error rates sufficiently to produce interior positive prices, as in Proposition 2. In addition, the resulting post-investment revenue should be greater than the incremental investment cost.
@@ 5 Welfare Analysis
Strictly profitable firm investments have a negative impact on consumer welfare and a strictly positive effect on firm welfare. The overall effect on welfare is positive (i.e., firm welfare + consumer welfare >> 0).
@@
$\text{Welfare}=\int_{0}^{\tilde{\alpha}}c_{i}(1)\,d\alpha+\int_{\tilde{\alpha}}^{1}c_{i}(2)\,d\alpha,$
where
@@
Firm investments on their inferior dimension have a strictly positive effect on consumers but a negative effect on overall welfare.
For example, an investment by firm 1 towards lowering its FP rate (its inferior dimension) would result in a strictly lower profit as shown by Theorem 4. Furthermore, firm 2 would experience a relatively larger drop in profit, due to losing market share but also having to reduce its price. However, consumers would be strictly better off due to lower prices. An interesting consequence of this is that firms can be incentivized to invest on their inferior dimension, which would be equivalent to reducing product differentiation in the market.
@@ 6 Summary and Conclusions
This paper develops a consumer-firm duopoly model to analyze optimal firm behavior through specialized investments. Each firm aims to minimize its model’s error, but this choice can often be suboptimal when firms compete and consumers have heterogeneous preferences. By decomposing model error into false positive and false negative rates, we allow firms to reduce error on one dimension at a time through investments.
@@
Our model assumes all information about consumer types is public information; introducing an element of private information into the consumer types could be interesting. Each consumer would be endowed with their own personal FP/FN rates and only buy a product from firms if the transaction is profitable error-wise. We assume that every consumer must not only purchase from a firm, but we can also deterministically pinpoint from which firm they will buy from. Implicitly, we assume a mature and entirely covered market in which we not only have a fixed market size with no growth but also enforce that all firms in the market are actually selling a product. A reasonable relaxation of the model could consider the possibility of shut-downs when their product is no longer profitable, and similarly the potential for market entry if a firm comes up with an innovative product.
@@ A.1 Equilibrium Prices in Perfectly Positively Correlated Model: $\alpha=\beta$
[…]
@@ A.2 Equilibrium Prices with Split Domination in Perfectly Negatively Correlated Model: $\alpha=1-\beta$
[…]
@@
The firms’ reaction functions can then be solved using casework on $\tilde{\alpha}$ and relevant FOCs. Note that if $\tilde{\alpha}=0$, then $R_{1}=0,R_{2}=p_{2}$. Since firms are not subject to direct costs in this model, firm 2 will set price $p_{2}=\bar{p}$ and $p_{1}=0$. Similarly, if $\tilde{\alpha}=1$, firm 2’s revenue will be 0 and equilibrium prices will be $p_{1}=\bar{p}$ and $p_{2}=0$. When $0<\tilde{\alpha}<1$, the reaction functions can be calculated by taking the FOCs of firm revenues with respect to prices.
@@ A.3 Equilibrium Prices with Strict Domination in Perfectly Negatively Correlated Model: $\alpha=1-\beta$
[…]
@@ B.2 Theorem 2
B.3 Corollary 1
We first explicitly derive an expression for $\frac{\partial R_{1}}{\partial FN_{1}}$ using a similar stream of logic to Theorem 2. In particular, let $a=FN_{2}-FN_{1}>0,b=FP_{1}-FP_{2}<0$ as in the proof for Theorem 1. Thus, from Claim 2 we have that prices are $p_{1}=\frac{2a+b}{3}$ and $p_{2}=\frac{a+2b}{3}$ and the general expression for revenue $R_{1}$ as:
$R_{1}=\tilde{\alpha}p_{1}=\left(\frac{p_{2}-p_{1}+a}{\Delta F_{1}-\Delta F_{2}}\right)\left(\frac{2a+b}{3}\right)$
We then want to compute $\frac{\partial R_{1}}{\partial FN_{1}}$:
$\frac{\partial R_{1}}{\partial FN_{1}}=$ $\frac{\partial}{\partial FN_{1}}(\tilde{\alpha})*p_{1}+\frac{\partial}{\partial FN_{1}}(p_{1})*\tilde{\alpha}$
$=$ $\left(-\frac{b}{3(\Delta F_{1}-\Delta F_{2})^{2}}\right)\left(\frac{2a+b}{3}\right)-\frac{2}{3}\left(\frac{2a+b}{3(\Delta F_{1}-\Delta F_{2})}\right)$
$=$ $\frac{(-1)\cdot(2a-b)\left[b+2(\Delta F_{1}-\Delta F_{2})\right]}{3(\Delta F_{1}-\Delta F_{2})^{2}}$
which is always negative as expected. Then, consider the standard firm profit maximization problem and the corresponding FOC:
$\Pi_{1}=R_{1}(p_{1},p_{2})-C_{1}(FP_{1},FN_{1},S_{1})$
$\frac{\partial\Pi_{1}}{\partial FN_{1}}=$ $\frac{\partial R_{1}}{\partial FN_{1}}-\frac{\partial C_{1}}{\partial FN_{1}}=0$
$\Rightarrow$ $\frac{(-1)\cdot(2a-b)\left[b+2(\Delta F_{1}-\Delta F_{2})\right]}{3(\Delta F_{1}-\Delta F_{2})^{2}}=g(S_{1})s_{\overline{FN_{1}}}(FN_{1}-\overline{FN_{1}})$
which gives the desired relation once we substitute back in the expressions for $a$ and $b$. The analysis for the optimal size $S_{1}$ is straightforward: since the size variable is latent and does not directly affect firm revenue, we simply differentiate the cost function with respect to $S_{1}$ and set it equal to zero.
Appendix C Additional Graphs
We present a visual interpretation of Claim 4 obtained through numerical simulation. We construct a grid space on $\zeta$ with increments of 0.01 and plot the corresponding price and revenue under equilibria. Note that both price and revenue are highest when consumers all have negatively correlated cost sensitivities.
Figure 5: Price and revenue of both firms across varying proportions of positively correlated consumers $\zeta$